OpenStudy (anonymous):

Solve 2|5x − 2| + 4 = 18 for x. Show your work to receive full credit.

OpenStudy (anonymous):

Get the absolute value by itself. 2|5x-2| + 4 = 18 2|5x-2| = 14 |5x-2| = 7 Then solve for the negative and the positive of the right side. 5x-2 = 7 5x = 9 x = 9/5 5x-2 = -7 5x = -5 x = -1 So, x = 9/5 and -1